# Do Large Code Models Understand Programming Concepts? A Black-box Approach

### v2

@@ Abstract
Large Language Models’ success on text generation has also made them better at code generation and coding tasks. While a lot of work has demonstrated their remarkable performance on tasks such as code completion and editing, it is still unclear as to why. We help bridge this gap by exploring to what degree auto-regressive models understand the logical constructs of the underlying programs. We propose Counterfactual Analysis for Programming Concept Predicates (CACP) as a counterfactual testing framework to evaluate whether Large Code Models understand programming concepts. With only black-box access to the model, we use CACP to evaluate ten popular Large Code Models for four different programming concepts. Our findings suggest that current models lack understanding of concepts such as data flow and control flow.
@@ 1 Introduction
Language Language Models (LLMs) have demonstrated remarkable performance on a variety of automated programming tasks, such as code completion (Austin et al., 2021; Fried et al., 2022), code repair (Jiang et al., 2021; Joshi et al., 2023), and code translation (Pan et al., 2023; Chen et al., 2023). Automating a programming task is a complex problem that requires understanding many concepts in the underlying code. These concepts include how variables are stored, accessed, and modified in memory; how execution proceeds across various constructs; and how different parts of the code compose sequentially or in parallel to perform a computation. We refer to these concepts as Programming-Concept Predicates (PCPs). Despite their remarkable performance, to what degree LLMs understand the PCPs in the programs they manipulate remains unclear.
@@
In this work, we consider the problem of evaluating a given model’s understanding of programming concepts. We focus on four PCPs that represent classical concepts in the programming analysis literature (Allen, 1970; Fosdick and Osterweil, 1976; Lin and Wu, 2008; Dart and Zobel, 1992):
@@ 2 Background and Related Work
Large language models (LLMs) have shown strong performance on a variety of code tasks, from code completion (Austin et al., 2021; Fried et al., 2022), to code translation (Pan et al., 2023; Chen et al., 2023), and to code repair (Jiang et al., 2021; Joshi et al., 2023). A code LLM takes as input a sequence of natural-language instructions and a sequence of code statements (i.e., a partial program) and outputs another partial program (depending on the task). We consider the general case where the task of interest has an associated function (called the attribution function) that determines whether the output of the model satisfies the input instruction. For generative tasks for code such as code completion or code repair, it is common to use program testing as attribution function, where the output program is executed against a test suite.
@@ 3.1 Notation
where $\mathcal{H}$ is the space of instructions and $\mathcal{X},\mathcal{Y}\in\mathcal{P}$ with $\mathcal{P}$ being the space of programs. For code completion, $\mathcal{H}$ is the docstring or the problem specification in natural language, and $\mathcal{X}$ and $\mathcal{Y}$ are program prefixes and completions, respectively. An attribution function $\mathsf{A}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\rightarrow\{0,1\}$ evaluates if the model output satisfies the instruction. For code completion, a common attribution function evaluates if the completed program passes the unit tests specified by the problem. Also, let $O_{h\times x}=\{y\;|\;y\in\mathcal{Y},\mathsf{A}(h,x,y)=1\}$ be the set of correct outputs for a given instruction-input pair, where $x\in\mathcal{X},h\in\mathcal{H}$.
@@ 3.2 Requirements
    Correctness: A counterfactual is considered correct if it achieves a desired outcome. For programs, this would mean that the perturbed program should still be able to solve the task described by the instructions. We use the task’s attribution function to verify this condition. Specifically, for a model $\mathsf{M}$, a counterfactual pair $x,x^{\prime}\in\mathcal{X}$, associated problem description $h\in\mathcal{H}$ and corresponding attribution function $\mathsf{A}$, we ensure that $|O_{h\times i}|>0\;\;\forall i\in\{x,x^{\prime}\}$.
@@
Figure 2: This figure illustrates the counterfactual generation pipeline of CACP. It consists of two stages. First, the reference solution for the problem is perturbed using predicate-specific mutations. Second, both the original and the perturbed solution are cut at the same location to generate a pair of counterfactual inputs.
@@ 3.3 Mutations for Counterfactual Programs
Now, we discuss how CACP generates counterfactual programs that satisfy the above requirements. CACP automates the CF generation process using mutations. These are transformation functions that perturb programs with respect to specific concepts, i.e., $\mathsf{T}_{\mathsf{p}_{k}}:\mathcal{X}\rightarrow\mathcal{X}$ where $\mathsf{p}_{k}$ is the target PCP. A PCP can have more than one associated mutation. Given an input program $x\in\mathcal{X}$, the mutation function is then used to generate a counterfactual $x_{\mathsf{p}_{k}}=\mathsf{T}_{\mathsf{p}_{k}}(x)\in\mathcal{X}$. Our comprehensive review of the program analysis literature revealed four themes of studied program predicates: control flow (Allen, 1970; Yang et al., 2015), data flow (Fosdick and Osterweil, 1976; Nilsson-Nyman et al., 2009), identifier names (Lin and Wu, 2008), and data types (Dart and Zobel, 1992; Allamanis et al., 2020). As we study weakly typed programs (for instance, Python), we consider four distinct PCPs that cover the first three themes. Next, we show how CACP automates the generation of these four distinct PCPs (also illustrated in Figure 2).
@@
Variable-Name Invariance: Next, we evaluate the PCP: Variable names do not affect the semantics of a program. Here, we generate counterfactuals by renaming variables. We consider two variants of this mutation — renaming to random strings and permuting or shuffling existing names between variables. For the first variant, we substitute variable names with randomly generated strings of five characters. For the second variant, we shuffle names among the variables defined in the program. This mutation satisfies – (1) Correctness: we ensure that the counterfactual is semantically equivalent by consistently substituting each variable; (2) Validity: We only substitute user-defined variables and do not rename reserved keywords; (3) Specificity: We do not substitute function parameters as they may be matched using names.
@@ 3.4 Measuring Counterfactual Effect
An Average Mutation Effect with a small magnitude indicates a better understanding of the PCP. On the other hand, a large magnitude indicates poor understanding since the model performs worse after the mutation. Note that this formulation is similar to the Average Treatment Effect used in counterfactual analysis (Pearl, 2009). The treatment Effect is defined for the output of the model, whereas we compute the Mutation Effect using the attribution function.
@@ 4.2 CACP Counterfactual Generation
Variable-Name Invariance: We only consider mutations where at least one variable occurrence is renamed in the initial 75% of the program. This ensures that counterfactual input is not identical to the original. We cut off the trailing 25% and use the rest as the counterfactuals.
@@ 4.3 CACP Effect Measurement
There are two primary approaches to evaluating the generations of a code-completion task—testing and exact string matching. Exact string-matching techniques like CodeBLEU (Ren et al., 2020) and chrF (Popović, 2015) evaluate generations by computing the distance from the reference solution. However, such match-based metrics are unable to account for the large space of programs that are functionally equivalent to, yet syntactically distinct from, a reference solution and thus underestimate the capabilities of a model that understands programming concepts. Testing provides a more direct evaluation, where a generation is deemed correct if it passes all the unit tests for that code-completion instance. Therefore, we use unit-test correctness as the attribution function for computing the Average Mutation Effect. We generate candidate solutions by querying the model on both the original input as well as the counterfactual. Then, we execute the candidate solutions against the test cases, resulting in one of two outcomes: passing all test cases or at least one failure. Note that we only consider problems where the model generates a successful completion for the original (non-perturbed) input, the perturbed input, or both. The cases where the model fails for both the original and perturbed inputs are not necessarily informative about the impact of the PCP, and we discard them. In that case, the perturbed inputs are not considered as counterfactual.
@@ 5 Experiments
We evaluate ten popular LLMs and show that they suffer significant drops in unit test correctness for mutations on Variable-Names, IfElse-Flip, and DefUse-Break, leading to Average Mutation Effects as high as $34\%$. The effect is smaller in magnitude for Independent-Swap. Overall, these results suggest that current models lack understanding of program predicates.
Q2: How does the Average Mutation Effect depend on LLM parameters?
@@
We analyze the correlation between pairs of mutations and show that all pairs exhibit low correlation apart from the two Variable Names mutations. In the case of StarCoder (Li et al., 2023), our analysis suggests a relation between AME for the IfElse-Flip mutation and the frequency of appearance of different relational operators in the model’s training data.
@@
Figure 3: Number of Instances where the unit test correctness of the counterfactual input changed / not changed for Starcoder. Independent-Swap: SWAP, IfElse-Flip: IFFP, Variable Names Random: RAND, Variable Names Shuffle: SHUF, DefUse Break: DFBR
Table 2: We compute the average mutation effect using the Pass/Fail attribute function as described in subsection 4.3. Accuracy is computed by averaging over 10 completions per problem for both original and perturbed partial programs. Also, we only consider problems where the model achieves non zero accuracy on the original setting.
Average Mutation Effect
@@ 5.1 Experimental Setup
Datasets and mutations. We instantiate CACP using three popular code generation benchmarks — HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and CodeContests (Li et al., 2022). All of the problems in these datasets include a reference solution, which is used to generate counterfactual pairs as described in section 4. Since not every mutation applies to all reference solutions, the final number of counterfactual pairs differs based on the mutation type. As shown in Table 1, mutations related to Variable Names can be applied to almost all solutions, whereas mutations related to control-flow or def-use are more selective. In this evaluation, we focus on Python programs, but our methodology applies to any programming language. We use libCST for parsing and manipulating source code for our mutations.
Models. We use CACP to evaluate popular models, including Llama 2 (Touvron et al., 2023) and PaLM (Anil et al., 2023). We also evaluate counterparts of these models that are fine-tuned for coding tasks – Code Llama (Roziere et al., 2023) and PaLM 2-$S^{*}$ (Anil et al., 2023). Finally, we also evaluate the popular open source code LLM Starcoder (Li et al., 2023). We set the temperature to 0 for all models to have deterministic results.
Figure 4: Average Mutation Effect as a function of model size (number of parameters in Billions). The different model classes are depicted using different colors.
Figure 5: Correlation between Average Mutation Effect values across pairs of mutations. The number of samples used to compute each value depends on the size of the intersection of the two mutation types. Independent-Swap: SWAP, IfElse-Flip: IFFP, Variable Names Random: RAND, Variable Names Shuffle: SHUF
@@ 5.2 Average Mutation Effect
Table 2 shows the Average Mutation Effect (AME) for the three datasets, five mutations, and ten models. The table shows that the original unit test correctness rates vary across models. AME values are non-zero, which suggests that models do not fully understand the evaluated PCPs. In the case of the Variable-Names and IfElse-Flip perturbations, AME values are as high as $33\%$. On the other hand, the Independent-Swap mutation is the most well-understood. While most mutations have similar effects across the two kinds of datasets, the DefUse-Break perturbation shows a relatively lower effect on the HumanEval and MBPP datasets. This is likely due to the small number of valid problems — only 22.
Across Models: For Variable-Name related perturbations, we can observe that smaller models perform worse and larger models do better. This is evident in Figure 4, which shows the AME as a function of the model size. Secondly, models trained on code (StarCoder) or fine-tuned on code (Llama Code, PaLM 2-$S^{*}$) perform better than models that are not. Perturbations related to control flow and data flow follow a similar trend for model size, but code fine-tuning does not always seem to improve performance.

### v3

# Do Large Code Models Understand Programming Concepts? Counterfactual Analysis for Code Predicates

@@ Abstract
Large Language Models’ success in text generation has also made them better at code generation and coding tasks. While a lot of work has demonstrated their remarkable performance on tasks such as code completion and editing, it is still unclear as to why. We help bridge this gap by exploring to what degree auto-regressive models understand the logical constructs of the underlying programs. We propose Counterfactual Analysis for Programming Concept Predicates (CACP) as a counterfactual testing framework to evaluate whether Large Code Models understand programming concepts. With only black-box access to the model, we use CACP to evaluate ten popular Large Code Models for four different programming concepts. Our findings suggest that current models lack understanding of concepts such as data flow and control flow.
Machine Learning, ICML
@@ 1 Introduction
Language Language Models (LLMs) have demonstrated remarkable performance on a variety of automated programming tasks, such as code completion (Austin et al., 2021; Fried et al., 2022), code repair (Jiang et al., 2021; Joshi et al., 2023), and code translation (Pan et al., 2023; Chen et al., 2023b). Automating a programming task is a complex problem that requires understanding many concepts in the underlying code. These concepts include how variables are stored, accessed, and modified in memory; how execution proceeds across various constructs; and how different parts of the code compose sequentially or in parallel to perform a computation. We refer to these concepts as Programming-Concept Predicates (PCPs). Despite their remarkable performance, the degree to which LLMs understand the PCPs in the programs they manipulate remains unclear.
@@
In this work, we consider the problem of evaluating a given model’s understanding of programming concepts. We focus on four PCPs that represent classical concepts in the program analysis literature (Allen, 1970; Fosdick & Osterweil, 1976; Lin & Wu, 2008; Dart & Zobel, 1992):
@@ 2 Background and Related Work
Large language models (LLMs) have shown strong performance on a variety of code tasks, from code completion (Austin et al., 2021; Fried et al., 2022), to code translation (Pan et al., 2023; Chen et al., 2023b), and to code repair (Jiang et al., 2021; Joshi et al., 2023). A code LLM takes as input a sequence of natural-language instructions and a sequence of code statements (i.e., a partial program) and outputs another partial program (depending on the task). We consider the general case where the task of interest has an associated function (called the attribution function) that determines whether the output of the model satisfies the input instruction. For generative tasks for code such as code completion or code repair, it is common to use program testing as attribution function, where the output program is executed against a test suite.
@@
Robustness of Code Models.
Recent work has studied the robustness of code models against both natural and adversarial perturbations. Shirafuji et al. (2023) & Wang et al. (2022) study robustness of code completion against different representations for the problem description as well as the input program. Henkel et al. (2022); Jha & Reddy (2023) demonstrate that function name prediction models can be attacked using semantic preserving transformation applied to the input program. Chen et al. (2023a) have similar findings for the code summarization task. In this work, we focus on evaluating the understanding of specific programming concepts. Our approach is based on causal analysis, which involves carefully designing counterfactuals and attributing their effect. Similar to robustness, our approach also involves mutating programs and performing inference. However, our mutations are aimed at generating counterfactuals and need to ensure that the input for the original prompt and the counterfactual prompt differ only along the concept to be tested.
@@ 3.1 Notation
where $\mathcal{H}$ is the space of instructions and $\mathcal{X},\mathcal{Y}\in\mathcal{P}$ with $\mathcal{P}$ being the space of programs. For code completion, $\mathcal{H}$ is the docstring or the problem specification in natural language, and $\mathcal{X}$ and $\mathcal{Y}$ are program prefixes and completions, respectively. An attribution function $\mathsf{A}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\rightarrow\{0,1\}$ evaluates if the model output satisfies the instruction. Also, let $O_{h\times x}=\{y\;|\;y\in\mathcal{Y},\mathsf{A}(h,x,y)=1\}$ be the set of correct outputs for a given instruction-input pair, where $x\in\mathcal{X},h\in\mathcal{H}$. For code completion, a common attribution function evaluates if the completed program passes the unit tests specified by the problem.
@@ 3.2 Requirements
    Correctness: A counterfactual needs to correctly solve the original task. For programs, this would mean that the perturbed program should still be able to solve the task described by the instructions. We use the task’s attribution function to verify this condition. Specifically, for a model $\mathsf{M}$, a counterfactual pair $x,x^{\prime}\in\mathcal{X}$, associated problem description $h\in\mathcal{H}$ and corresponding attribution function $\mathsf{A}$, we ensure that $|O_{h\times i}|>0\;\;\forall i\in\{x,x^{\prime}\}$.
@@
Figure 2: The counterfactual generation pipeline of CACP. It consists of two stages. First, the reference solution for the problem is perturbed using predicate-specific mutations. Second, both the original and the perturbed solution are cut at the same location to generate a pair of counterfactual inputs.
@@ 3.3 Mutations for Counterfactual Programs
Now, we discuss how CACP generates counterfactual programs that satisfy the above requirements. CACP automates the CF generation process using mutations. These are transformation functions that perturb programs with respect to specific concepts, i.e., $\mathsf{T}_{\mathsf{p}_{k}}:\mathcal{X}\rightarrow\mathcal{X}$ where $\mathsf{p}_{k}$ is the target PCP. A PCP can have more than one associated mutation. Given an input program $x\in\mathcal{X}$, the mutation function is then used to generate a counterfactual $x_{\mathsf{p}_{k}}=\mathsf{T}_{\mathsf{p}_{k}}(x)\in\mathcal{X}$. Our comprehensive review of the program analysis literature revealed four themes of studied program predicates: control flow (Allen, 1970; Yang et al., 2015), data flow (Fosdick & Osterweil, 1976; Nilsson-Nyman et al., 2009), identifier names (Lin & Wu, 2008), and data types (Dart & Zobel, 1992; Allamanis et al., 2020). As we study weakly typed programs (for instance, Python), we consider four distinct PCPs that cover the first three themes. Next, we show how CACP automates the generation of these four distinct PCPs (also illustrated in Figure 2).
@@
Variable-Name Invariance: Next, we evaluate the PCP: Variable names do not affect the semantics of a program. Here, we generate counterfactuals by renaming variables. We consider two variants of this mutation — renaming to random strings and permuting or shuffling existing names between variables. For the first variant, we substitute variable names with randomly generated strings of five characters. For the second variant, we shuffle names among the variables defined in the program. This mutation satisfies – (1) Correctness: we ensure that the counterfactual is semantically equivalent by consistently substituting each variable; (2) Validity: We only substitute user-defined variables and do not rename reserved keywords; (3) Specificity: We do not substitute function parameters as their names decide the order in which arguments are passed during invocation.
@@ 3.4 Measuring Counterfactual Effect
$\mathsf{AME}$ with a small magnitude indicates a better understanding of the PCP. On the other hand, a large magnitude indicates poor understanding since the model performs worse after the mutation. Note that this formulation is similar to the Average Treatment Effect used in counterfactual analysis (Pearl, 2009). The treatment Effect is defined for the output of the model, whereas we compute the Mutation Effect using the attribution function.
@@ 4.2 CACP Counterfactual Generation
Variable-Name Invariance: We only consider mutations where at least one variable appearance is renamed in the initial 75% of the program. This ensures that counterfactual input is not identical to the original. We cut off the trailing 25% and use the rest as the counterfactuals.
@@ 4.3 CACP Effect Measurement
There are two primary approaches to evaluating the generations of a code-completion task—testing and exact string matching. Exact string-matching techniques like CodeBLEU (Ren et al., 2020) and chrF (Popović, 2015) evaluate generations by computing the distance from the reference solution. However, such match-based metrics are unable to account for the large space of programs that are functionally equivalent to, yet syntactically distinct from, a reference solution and thus underestimate the capabilities of a model that understands programming concepts. Testing provides a more direct evaluation, where a generation is deemed correct if it passes all the unit tests for that code-completion instance. Therefore, we use unit-test correctness as the attribution function for computing the $\mathsf{AME}$. We generate candidate solutions by querying the model on both the original input as well as the counterfactual. Then, we execute the candidate solutions against the test cases, resulting in one of two outcomes: passing all test cases or at least one failure. Note that we only consider problems where the model generates a successful completion (i.e. passing all test cases) for the original (non-perturbed) input, the perturbed input, or both. The cases where the model fails for both the original and perturbed inputs are not necessarily informative about the impact of the PCP, and we discard them. In that case, the perturbed inputs are not considered as counterfactual.
@@ 5 Experiments
We evaluate ten popular LLMs and show that they suffer significant drops in unit test correctness for mutations on Variable-Names, IfElse-Flip, and DefUse-Break, leading to $\mathsf{AMEs}$ as high as $34\%$. The effect is smaller in magnitude for Independent-Swap. Overall, these results suggest that current models lack understanding of program predicates.
Q2: How does the Average Mutation Effect ($\mathsf{AME}$) depend on LLM parameters?
@@
We analyze the correlation between pairs of mutations and show that all pairs exhibit low correlation apart from the two Variable Names mutations. In the case of StarCoder (Li et al., 2023), our analysis suggests a relation between $\mathsf{AME}$ for the IfElse-Flip mutation and the frequency of appearance of different relational operators in the model’s training data.
@@
Table 2: We compute the $\mathsf{AME}$ using the Pass/Fail attribute function as described in subsection 4.3. We only consider problems where the model achieves non zero accuracy on either the original or the counterfactual setting.
Average Mutation Effect (AME)
@@ 5.1 Experimental Setup
Datasets and mutations. We instantiate CACP using three popular code generation benchmarks — HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and CodeContests (Li et al., 2022). All of the problems in these datasets include a reference solution, which is used to generate counterfactual pairs as described in section 4. Since not every mutation applies to all reference solutions, the final number of counterfactual pairs differs based on the mutation type. As shown in Table 1, mutations related to Variable Names can be applied to almost all solutions, whereas mutations related to control-flow or def-use are more selective. In this evaluation, we focus on Python programs, but our methodology applies to any programming language. We use libCST (LibCST Developers, ) for parsing and manipulating source code for our mutations.
Models. We use CACP to evaluate popular models, including Llama 2 (Touvron et al., 2023) and PaLM (Anil et al., 2023). We also evaluate counterparts of these models that are fine-tuned for coding tasks – Code Llama (Roziere et al., 2023) and PaLM 2-$S^{*}$ (Anil et al., 2023). Finally, we also evaluate the popular open source code LLM StarCoder (Li et al., 2023). We set the sampling temperature to 0 for all models to have deterministic results.
Figure 3: $\mathsf{AME}$ as a function of model size (number of parameters in Billions). The different model classes are depicted using different colors.
Figure 4: Correlation between $\mathsf{AME}$ values across pairs of mutations. The number of samples used to compute each value depends on the size of the intersection of the two mutation types. Independent-Swap: SWAP, IfElse-Flip: IFFP, Variable Names Random: RAND, Variable Names Shuffle: SHUF
@@ 5.2 Average Mutation Effect
Table 2 shows the $\mathsf{AME}$ for the three datasets, five mutations, and ten models. The table shows that the original unit test correctness rates vary across models. $\mathsf{AME}$ values are non-zero, which suggests that models do not fully understand the evaluated PCPs. In the case of the Variable-Names and IfElse-Flip perturbations, $\mathsf{AME}$ values are as high as $33\%$. On the other hand, the Independent-Swap mutation is the most well-understood. While most mutations have similar effects across the two kinds of datasets, the DefUse-Break perturbation shows a relatively lower effect on the HumanEval and MBPP datasets. This is likely due to the small number of valid problems — only 22.
Across Models: For Variable-Name related perturbations, we first observe that smaller models perform worse and larger models do better. This is evident in Figure 3, which shows the $\mathsf{AME}$ as a function of the model size. Secondly, models trained on code (StarCoder) or fine-tuned on code (Llama Code, PaLM 2-$S^{*}$) perform better than models that are not. Perturbations related to control flow and data flow follow a similar trend for model size, but code fine-tuning does not always seem to improve performance. GPT4 performs much better on HumanEval and MBPP, but is similar to the other models for the CodeContests dataset.
@@
Effect of cutoff point: In subsection 4.2, we describe that we keep 75% of the program as the prefix for generating counterfactuals for the Independent-Swap, DefUse-Break and Variable Name Invariance mutations. To study the effect of the cutoff point, we evaluated counterfactuals generated using the same set of programs but cut at different places. In Table 4, we present the original accuracy and the $\mathsf{AME}$ for the Starcoder model. This does not include the IfElse-Flip mutation since in that case the cut depends on the location of the If block. We find that an earlier cut leads to a decrease in the original accuracy as well as a higher $\mathsf{AME}$. This is expected since cutting earlier increases the complexity of the completion task. However, we observe that $\mathsf{AME}$ is relatively more stable than the original accuracy. This suggests that $\mathsf{AME}$ is a good measure of the model’s understanding, irrespective of the complexity of the coding task.
Code Repair Task: We also evaluate CACP for the code repair task. We use the HumanEvalPack (Muennighoff et al., 2023) dataset which is an extension of HumanEval to also include the Code Repair task. This dataset is constructed by manually adding a bug to each solution in HumanEval. For this task, the model is tasked with fixing the bug and generating the correct solution. In this case, we generate counterfactuals by applying mutations on the buggy solution. In Table 5, we show the performance of Octocoder (Muennighoff et al., 2023) which is an instruction-tuned version of Starcoder. Similar to code completion, we observe a high average mutation effect which suggests a lack of understanding for the Code Repair task as well.
@@ 6 Future Work
Table 4: $\mathsf{AME}$ for different cutoff settings when evaluating Starcoder. A lower prefix ratio implies an earlier cut. Independent-Swap: SWAP, Variable Names Random: RAND, Variable Names Shuffle: SHUF, DefUse-Break: DUBR
Prefix
Ratio
Orig.
Acc.
RAND SHUF SWAP DUBR
0.4 - 0.6 32.8 % 18.7 % 21.4 % 08.0 % 26.0 %
0.6 - 0.8 50.1 % 16.7 % 19.6 % 25.0 % 06.0 %
0.8 - 1.0 60.1 % 12.4 % 17.7 % 19.6 % 04.0 %
@@
Table 5: $\mathsf{AME}$ for the code repair task. We evaluate OctoCoder (Muennighoff et al., 2023) on countefactuals generated on the code repair benchmark from HumanEvalPack. Independent-Swap: SWAP, Variable Names Shuffle: SHUF, IfElse-Flip: IFFP
Original Accuracy SHUF IFFP SWAP
15 % 36 % 33 % 15 %